# Time domain astrophysics with transient sources

### v1

@@
(Accepted with minor review on 29 May 2025)
@@
The developed techniques provides a significant computational advantage and are useful analyzing of data characterized by low-count statistics (i.e., low photon count rates in c/s), as they allow overcoming the limitations associated with traditional fixed bin-size methods.
@@ 1 Introduction
GRBs are short, intense, and unrepeatable flashes of radiation, with a spectral energy distribution peaking in the gamma-ray band (D’Avanzo 2015). The theoretical isotropic energy released can reach up to $10^{55}$ erg (Wu et al. 2012; Dado & Dar 2022), over a period ranging from fractions of a second up to several thousand seconds (von Kienlin et al. 2020). However, jet-like emission in GRBs can reduce the required energy budget to produce the observed brightness by at least a factor of 100, as the energy is directed narrowly rather than spread isotopically (Sari et al. 1999). The GRB generation is associated with the gravitational collapse of a massive star (Piran 1999; Campana et al. 2008) or the coalescence of two neutron stars in an extremely close binary system because of the emission of gravitation waves. Indeed the GRB 170817A event (Savchenko et al. 2017; Goldstein et al. 2017) is believed to be an example of the latter option. Another hypothesis proposes that the merger of a black hole and a neutron star could serve as a GRB progenitor. However, no direct observational evidence currently supports this possibility (Mochkovitch et al. 1993; Gompertz et al. 2020).
@@ 2.2 How to built a light curve
If $N$ photons are detected in a certain $\rm{\Delta t}$, the mode of the Poisson distribution, $\rm{r}={\rm{N}}/{\rm{\Delta t}}$, is the most probable rate value in such interval. This value is an average rate associated with a confidence interval which depends on the chosen CL and the number of photons $N$, as as illustrated in Figure 2.
@@
[…]
1 $0.29$ $2.49$ 00 $2$
5 $0.61$ $1.52$ $0.55$ $1.45$
10 $0.72$ $1.35$ $0.68$ $1.32$
50 $0.86$ $1.15$ $0.86$ $1.14$
100 $0.90$ $1.10$ $0.90$ $1.10$
Table 1: Comparison between the [$\rm\epsilon^{-}(N,CL)$ , $\rm\epsilon^{+}(N,CL)$] confidence interval and the $1\sigma$ Gaussian relative confidence interval $1\pm\frac{1}{\sqrt{\rm{N}}}$ as function of the number of counts $N$.
@@ 4 Errors treatment
Due to the probabilistic nature of the process, when two identical detectors observe the same GRB, the obtained rates, $r_{1}$ and $r_{2}$, differ due to Poisson fluctuations. Consequently, the delay estimate $\tau$ generally differs from the expected value $\rm{\tau}=0\text{s}$ when cross-correlating light curves of the same event observed by two detectors positioned side by side. For instance, we consider the ToA lists associated with the GRB 0908207 as observed by two Fermi/GBM detectors (NaI detector 1 and NaI detector 5, 10 keV - 900 keV energy band). These are physically separated by a maximum distance of 5 m, corresponding to the diagonal of the almost cubic shape of the Fermi satellite, which implies a maximum theoretical delay of $\rm{\tau}_{th}=15\rm{ns}$ (Bissaldi et al. 2009; Meegan et al. 2009). We must also consider that the two detectors had different pointing directions at the time of the burst. Consequently, the observed photon count and respective rate vary depending on the off-axis angle relative to the source.
@@
Taken at face value, this result would imply that some unknown systematic effect has biased the measurement. The single CCF formally yields a significant lag with a minimal uncertainty, as it inherently captures the particular statistical fluctuation in the pair of detector measurements. This small uncertainty, however, is purely mathematical and pertains only to the statistical variation specific to that individual realization. Repeating the measurement under identical conditions would result in a different lag estimate due to random fluctuations. Therefore, a Monte Carlo (MC) simulation approach is necessary to accurately estimate the overall uncertainty, incorporating the full range of possible statistical fluctuations (Zhang et al. 1999). The MC distribution of delays is centered around the best experimental estimate $\rm{\mu}$ and the associated error is the standard deviation of the distribution $\sigma$.
@@ 4.1 The methods
where RND(0,1) is a Poisson flat distribution that generates a random number between 0 and 1. ${\rm{T\_SIM[N]}}$ is the ToA to be recursively simulated, starting with a previous time $\rm T\_SIM[N-1]$. In the first step, $\rm T\_SIM[N-1]$ is $\rm{t}_{1}$. This approach emulates the detector measurement process, using a Poisson arrival process applied to the rate of the observed signal.
@@
By generating a large number of couples of ToA, each belonging to one Pool, and cross-correlating the light curves obtained from them, we obtain a large number of delays with an approximately Gaussian distribution because of the center limit theorem. The mean of this distribution represents the expected delay, while the sigma indicates the uncertainty.
@@ 5.1 MDP and DP methods comparison
Figure 8: DP and MDP methods comparison. The plots on the left show the experimental delays estimated via MC procedures and the associated error as the vertical error bar. The plots on the right show the residual distribution for each method in units of sigma.
@@ 6 Conclusion
We can therefore confirm the capability of the MDP to deliver reliable scientific results, providing a significant increase in both accuracy and computational efficiency. Furthermore, the developed techniques demonstrate crucial effectiveness in low-statistics regimes, where traditional methods may struggle to yield consistent results. These techniques remain effective regardless of the transient signal luminosity, although the precision of the estimated lag improves with the increasing number of source-associated photons.
@@ Acknowledgements.
This article was produced while attending the PhD program in PhD in Space Science and Technology at the University of Trento, Cycle XXXVIII. A.T. acknowledges financial support from “ASI-INAF Accordo Attuativo HERMES Pathinder operazioni n. 2022-25-HH.0” and the basic funding program of the Ioffe Institute FFUG-2024-0002.

### v2

@@
(Accepted, 25 June 2025)
@@
The developed techniques provide a significant computational advantage and are useful analysis of data characterized by low-count statistics (i.e., low photon count rates in c/s), as they allow overcoming the limitations associated with traditional fixed bin-size methods.
@@ 1 Introduction
GRBs are short, intense, and unrepeatable flashes of radiation, with a spectral energy distribution peaking in the gamma-ray band (D’Avanzo 2015). The theoretical isotropic energy released can reach up to $10^{55}$ erg (Wu et al. 2012; Dado & Dar 2022), over a period ranging from fractions of a second up to several thousand seconds (von Kienlin et al. 2020). However, jet-like emission in GRBs can reduce the required energy budget to produce the observed brightness by at least a factor of 100, as the energy is directed narrowly rather than spread isotopically (Sari et al. 1999). The GRB generation is associated with the gravitational collapse of a massive star (Piran 1999; Campana et al. 2008) or the coalescence of two neutron stars in an extremely close binary system because of the emission of gravitational waves. Indeed the GRB 170817A event (Savchenko et al. 2017; Goldstein et al. 2017) is believed to be an example of the latter option. Another hypothesis proposes that the merger of a black hole and a neutron star could serve as a GRB progenitor. However, no direct observational evidence currently supports this possibility (Mochkovitch et al. 1993; Gompertz et al. 2020).
@@ 2.2 How to built a light curve
If $N$ photons are detected in a certain $\rm{\Delta t}$, the mode of the Poisson distribution, $\rm{r}={\rm{N}}/{\rm{\Delta t}}$, is the most probable rate value in such interval. This value is an average rate associated with a confidence interval which depends on the chosen CL and the number of photons $N$, as illustrated in Figure 2.
@@
Table 1: Poisson vs. Gaussian $1\sigma$ relative confidence intervals
[…]
1 0.29 2.49 0 2
5 0.61 1.52 0.55 1.45
10 0.72 1.35 0.68 1.32
50 0.86 1.15 0.86 1.14
100 0.90 1.10 0.90 1.10
 [^1] Comparison between the Poisson confidence interval [$\epsilon^{-}(N,1\sigma),\,\epsilon^{+}(N,1\sigma)$] and the Gaussian approximation $1\pm 1/\sqrt{N}$ as a function of the number of counts $N$.
@@ 4 Errors treatment
Due to the probabilistic nature of the process, when two identical detectors observe the same GRB, the obtained rates, $r_{1}$ and $r_{2}$, differ due to Poisson fluctuations. Consequently, the delay estimate $\tau$ generally differs from the expected value $\rm{\tau}=0\text{s}$ when cross-correlating light curves of the same event observed by two detectors positioned side by side.
For instance, we consider the ToA lists associated with the GRB 0908207 (see Figure 4) as observed by two Fermi/GBM detectors (NaI detector 1 and NaI detector 5, 10 keV - 900 keV energy band). These are physically separated by a maximum distance of 5 m, corresponding to the diagonal of the almost cubic shape of the Fermi satellite, which implies a maximum theoretical delay of $\rm{\tau}_{th}=15\rm{ns}$ (Bissaldi et al. 2009; Meegan et al. 2009). We must also consider that the two detectors had different pointing directions at the time of the burst. Consequently, the observed photon count and respective rate vary depending on the off-axis angle relative to the source.
@@
Taken at face value, this result would imply that some unknown systematic effect has biased the measurement. The single CCF formally yields a significant lag with a minimal uncertainty, as it inherently captures the particular statistical fluctuation in the pair of detector measurements. This small uncertainty, however, is purely mathematical and pertains only to the statistical variation specific to that individual realization. Repeating the measurement under identical conditions would result in a different lag estimate due to random fluctuations.
Therefore, a Monte Carlo (MC) simulation approach is necessary to accurately estimate the overall uncertainty, incorporating the full range of possible statistical fluctuations (Zhang et al. 1999). The MC distribution of delays is centered around the best experimental estimate $\rm{\mu}$ and the associated error is the standard deviation of the distribution $\sigma$.
@@ 4.1 The methods
where RND(0,1) denotes a value drawn from a uniform distribution in the interval between 0 and 1. ${\rm{T\_SIM[N]}}$ is the ToA to be recursively simulated, starting with a previous time $\rm T\_SIM[N-1]$. In the first step, $\rm T\_SIM[N-1]$ is $\rm{t}_{1}$. This approach emulates the detector measurement process, using a Poisson arrival process applied to the rate of the observed signal.
@@
By generating a large number of couples of ToA, each belonging to one Pool, and cross-correlating the light curves obtained from them, we obtain a large number of delays with an approximately Gaussian distribution because of the central limit theorem. The mean of this distribution represents the expected delay, while the sigma indicates the uncertainty.
@@ 5.1 MDP and DP methods comparison
We note that the MDP method effectively mitigates intrinsic Poisson fluctuations in the input templates used by standard flux randomization methods or the DP method. These fluctuations would otherwise propagate and amplify across all Monte Carlo realizations, with a stronger effect as the chosen number of photons per bin is lower. While the method requires halving the available data at each step, resulting in an average loss of precision of approximately $\sqrt{2}$, it prevents the artificial amplification of Poisson noise. Consequently, it removes eventual bias in delay estimation introduced by Poissonian fluctuations in the original templates.
@@ 5.1.1 Proving that MDP method is fully independent
Figure 10: DP and MDP methods comparison. The plots on the left (DP and MDP) show the experimental delays estimated via MC procedures and the associated error as the vertical error bar. The plots on the right show the residual distribution for each method in units of sigma.
@@ 6 Conclusion
We can therefore confirm the capability of the MDP to deliver reliable scientific results, providing a significant increase in both accuracy and computational efficiency. We conclude that the MDP method reduces the impact of intrinsic Poisson fluctuations in input templates, avoiding their amplification in Monte Carlo simulations. Despite a $\sqrt{2}$ loss in precision due to halved statistics, it ensures unbiased delay estimates.
Furthermore, the developed techniques demonstrate crucial effectiveness in low-statistics regimes, where traditional methods may struggle to yield consistent results. These techniques remain effective regardless of the transient signal luminosity, although the precision of the estimated lag improves with the increasing number of source-associated photons.
@@ Acknowledgements.
This article was produced while attending the PhD program in PhD fin Space Science and Technology at the University of Trento, Cycle XXXVIII. A.T. acknowledges financial support from “ASI-INAF Accordo Attuativo HERMES Pathinder operazioni n. 2022-25-HH.0” and the basic funding program of the Ioffe Institute FFUG-2024-0002. We sincerely thanks Cristiano Guidorzi for his valuable comments and discussions.
@@ Appendix D Generalized inversion method analytical solution
Appendix E CCF examples
Figure 11 from page 12 to 16 shows the GRB considered in the comparison in Figure 10 between DP and MDP method.